Prove the following identity
Csc x - 2 sin x= cos (2x)/sin x

Answer by stanbon(75887)

Csc x - 2 sin x= cos (2x)/sin x
----
Multiply thru by sin(x) to get:
1 - 2sin^2(x) = cos(2xs)
------
That is one of the Pythagorean identities:
cos(2x) = cos(2x)
---
Cheers,
Stan H.